OpenStudy (anonymous):

A bag contains 10 white balls and 15 black balls. Two balls are drawn in succession .what is probability that one of them is black and the other is white.

OpenStudy (mathmate):

Given 10W 15B => total=25 Interpreting "drawn in succession" means no replacement. Assume balls are well mixed and selection is random. Need P(WB)+P(BW)=probability of drawing a black and a white, in any order. White drawn first. P(WB)=(10/25)*(15/24)=150/600=1/4 (25=original total, 24=no. of balls left after first draw). Go along the same lines for Black drawn first P(BW), and add together the two to get the final answer. Btw, this is a typical mathematics question, so please post in the math section in the future. You will probably get a faster response.

OpenStudy (anonymous):

Thanks a lot.☺️ but I posted it in math section and I don't get a reply so I post it in computer science.

OpenStudy (mathmate):

That's ok. What I would do is to first post in the appropriate section, and then as required, post in any other section with a link to the original question. This way, all responses are centralized in one place, even if there are 5 different posts for the same question.